### v1

@@ Abstract
We investigate the use of wavelet-space feature decomposition in neural super-resolution for rendering pipelines. Building on the DFASR framework, we introduce a wavelet-domain representation that separates low- and high-frequency details before reconstruction, enabling the network to better preserve fine textures while maintaining structural consistency. Unlike RGB-space regression, our approach leverages the stationary wavelet transform (SWT) to avoid spatial down-sampling, ensuring alignment across subbands and preserving shift invariance. The model predicts wavelet coefficients conditioned on spatial G-buffers and temporally warped history frames, which are then recombined through inverse wavelet synthesis. We conduct a comprehensive ablation study across wavelet families, transform types, and architectural variants, showing that incorporating SWT improves PSNR by up to 1.5 dB and reduces LPIPS by 17% on average, at a computational overhead of roughly +24 ms compared to out DFASR baseline. While absolute runtimes on our RTX 3050 mobile GPU are higher ( 141ms) than the original DFASR report on RTX 4090( 11ms), the relative overhead remains modest, suggesting that on higher-end GPUs our method would also remain real-time capable. Taken together, our results suggest that wavelet-domain representations are a principled and effective way to enhance perceptual quality in neural upscaling for graphics applications.
@@
Real-Time Rendering, Super-Resolution, Wavelet transform
@@ 1 Introduction
The demand for high-resolution graphics at interactive frame rates has pushed modern rendering pipelines to their computational limits. Physically based shading and real-time ray tracing improve visual fidelity but are often too expensive to run at native resolutions. Neural super-resolution has thus emerged as a practical alternative: frames are rendered at lower resolution and then reconstructed to the display resolution by learned models, balancing quality and efficiency.
@@
We propose a wavelet-domain formulation for neural super-resolution for real-time rendering pipeline. Instead of predicting pixel intensities, our model regresses wavelet coefficients that encode localized frequency information. This explicit spectral representation improves high-frequency reconstruction and reduces blur in textured regions. To avoid resolution loss across subbands, we adopt the stationary wavelet transform (SWT), with inverse reconstruction embedded in the network to allow end-to-end training. Our system supports arbitrary upscaling factors, similar to DFASR, but achieves sharper detail by operating in coefficient space. The approach provides an efficient path toward higher-fidelity rendering within real-time constraints.
@@ 2.1 Super-Resolution for Realtime-Rendering
Later work incorporates high-resolution G-buffers to provide geometric and material detail cheaply. FuseSR illustrates this trend, using a multi-resolution fusion architecture (“H-Net”) to align and fuse LR color with HR G-buffers, enabling temporally consistent $4\times$ and $8\times$ upsampling at 4K [9]. This mitigates LR information limits but still depends on discrete scales and accurate multi-resolution alignment.
@@ 3.1 Overview
Our method formulates real-time super-resolution as a wavelet-domain reconstruction problem. The goal is to recover a high-resolution frame $\hat{I}_{HR}$ from a combination of low-resolution shading, multi-resolution G-buffers, and temporal history. We denote the overall process as
@@ 3.2 Wavelet-Space Formulation
Figure 1: Single-level wavelet decomposition of an RGB image into LLLL, LHLH, HLHL, and HHHH subbands per channel. Our network predicts these subbands directly, then applies $\mathcal{W}^{-1}_{\tau,b}$ to reconstruct $\widehat{I}_{t}^{HR}$.
@@ 3.3 Data Preprocessing and Input Construction
with $\mathcal{M}$ denoting our wavelet-based reconstruction model. Here $G_{t}^{HR}$ is the high-resolution G-buffers, $Warp(\cdot)$ is temporal reprojection of the previous irradiance frame using motion vectors $MV_{t}$, and $\Phi=\{\Phi_{\text{spat}},\Phi_{\text{temp}}\}$ are spatial and temporal masks as intorduced in [7].
@@ 3.4 Network Architecture
Figure 3: Overview of our network architecture. The model consists of three feature extractors (for LR frames, HR G-buffers, and temporal inputs), a convolutional fusion branch, and an implicit neural representation (INR) branch with Fourier feature mapping. Their outputs are combined in the wavelet domain and reconstructed into the final image via inverse wavelet transform and pixel shuffle.
@@ 3.4.1 Feature Extractors
The LR-FE processes the pixel-unshuffled, demodulated low-resolution input frame $I_{t}^{LR}$, using a stack of $3\times 3$ convolutions to produce a feature map later routed to both fusion and INR pathways. The GB-FE encodes the high-resolution G-buffers $G_{t}^{HR}$ through residual blocks, extracting geometry- and material-aware information such as depth and surface orientation at display resolution. The T-FE receives the warped previous frame $\mathrm{Warp}(I_{t-1}^{HR},MV_{t})$ together with temporal masks, and is implemented with lightweight gated convolutions (LWGC) to balance motion sensitivity and efficiency.
@@ 3.4.5 Loss Functions
This formulation enforces accuracy across spectral, perceptual, and temporal dimensions, ensuring stable reconstructions under real-time rendering.
@@ 4.5 Runtime and Efficiency
Figure 6: Upscaling results on multiple scenes using NoWave, DFASR [7], and our SWT model. The first and second scenes show $2\times$ upscaling, while the third scene shows $3\times$ upscaling. Our SWT model consistently preserves edges and fine details better than the baselines, producing sharper and cleaner outputs across all tested scenarios.
@@ 5 Conclusion
Acknowledgments
The authors would like to express their sincere gratitude to Er. Saroj Shakya for his invaluable guidance and support as a co-supervisor throughout the duration of this project. We would also like to thank Ashim Tiwari for providing the necessary computational resources, which were essential for the successful completion of our work. Special thanks go to Ashim Bhharai for his practical assistance and hands-on support during the implementation phase of the project.

### v2

@@ Abstract
We investigate the use of wavelet-space feature decomposition in neural super-resolution for rendering pipelines. Building on neural upscaling frameworks, we introduce a wavelet-domain representation that separates low-frequency and high-frequency details before reconstruction, enabling the network to better preserve fine textures while maintaining structural consistency. Unlike RGB-space regression, our approach leverages the stationary wavelet transform (SWT) to avoid spatial down-sampling, ensuring alignment across subbands and preserving shift invariance. The model predicts wavelet coefficients conditioned on spatial G-buffers and temporally warped history frames, which are then recombined through inverse wavelet synthesis. We conduct a comprehensive ablation study across wavelet families, transform types, and architectural variants, showing that incorporating SWT improves PSNR by 1.5 dB and reduces LPIPS by 17% on average, with only a modest relative runtime overhead. Taken together, our results suggest that wavelet-domain representations a principled path toward higher-quality super-resolution in graphics applications.
@@
Computer graphics, Super-Resolution, Wavelet transform, Neural networks, Rendering pipelines
@@ 1 Introduction
The demand for high-resolution graphics at interactive frame rates has pushed modern rendering pipelines to their computational limits. Physically based shading and ray tracing improve visual fidelity but are often too expensive to run at native resolutions. Neural super-resolution has thus emerged as a practical alternative: frames are rendered at lower resolution and then reconstructed to the display resolution by learned models, balancing quality and efficiency.
@@
We propose a wavelet-domain formulation for neural super-resolution for rendering pipelines. Instead of predicting pixel intensities, our model regresses wavelet coefficients that encode localized frequency information. This explicit spectral representation improves high-frequency reconstruction and reduces blur in textured regions. To avoid resolution loss across subbands, we adopt the stationary wavelet transform (SWT), with inverse reconstruction embedded in the network to allow end-to-end training. Our system supports arbitrary upscaling factors, similar to DFASR, but achieves sharper detail by operating in coefficient space. The approach provides an efficient path toward higher-fidelity rendering.
@@ 2.1 Super-Resolution for Realtime-Rendering
Later work incorporates high-resolution G-buffers to provide geometric and material detail efficiently. FuseSR illustrates this trend, using a multi-resolution fusion architecture (“H-Net”) to align and fuse LR color with HR G-buffers, enabling temporally consistent $4\times$ and $8\times$ upsampling at 4K [9]. This mitigates LR information limits but still depends on discrete scales and accurate multi-resolution alignment.
@@ 3.1 Overview
Our method formulates super-resolution as a wavelet-domain reconstruction problem. The goal is to recover a high-resolution frame $\hat{I}_{HR}$ from a combination of low-resolution shading, multi-resolution G-buffers, and temporal history. We denote the overall process as
@@ 3.2 Wavelet-Space Formulation
Figure 1: Single-level wavelet decomposition of an RGB image into LLLL, LHLH, HHHH, and HLHL subbands per channel. Our network predicts these subbands directly, then applies $\mathcal{W}^{-1}_{\tau,b}$ to reconstruct $\widehat{I}_{t}^{HR}$.
@@ 3.3 Data Preprocessing and Input Construction
with $\mathcal{M}$ denoting our wavelet-based reconstruction model. Here $G_{t}^{HR}$ is the high-resolution G-buffers, $Warp(\cdot)$ is temporal reprojection of the previous irradiance frame using motion vectors $MV_{t}$, and $\Phi=\{\Phi_{\text{spat}},\Phi_{\text{temp}}\}$ are spatial and temporal masks as introduced in [7].
@@ 3.4 Network Architecture
Figure 3: Overview of our network architecture. The model consists of three feature extractors (for LR frames, HR G-buffers, and temporal inputs), a convolutional fusion branch, and an implicit neural representation (INR) branch with Fourier feature mapping. Their outputs are combined in the wavelet domain and reconstructed into the final image via inverse wavelet transform and pixel shuffle. Convolutional blocks are $3\times 3$ with padding 1, except the RIR-blocks which consist of $1\times 1$ expand/reduce layers with an inner $3\times 3$ convolution.
@@ 3.4.1 Feature Extractors
The LR-FE processes the pixel-unshuffled, demodulated low-resolution input frame $I_{t}^{LR}$, using a stack of $3\times 3$ convolutions to produce a feature map later routed to both fusion and INR pathways. The GB-FE encodes the high-resolution G-buffers $G_{t}^{HR}$ through RIR blocks, extracting geometry- and material-aware information such as depth and surface orientation at display resolution. The T-FE receives the warped previous frame $\mathrm{Warp}(I_{t-1}^{HR},MV_{t})$ together with temporal masks, and is implemented with lightweight gated convolutions (LWGC) to balance motion sensitivity and efficiency.
@@ 3.4.5 Loss Functions
This formulation enforces accuracy across spectral, perceptual, and temporal dimensions, ensuring stable reconstructions for rendering.
@@ 4.5 Runtime and Efficiency
Figure 6: Upscaling results on multiple scenes using NoWave, DFASR [7], and our SWT model. The first scene (Asian Village) demonstrates that our SWT model recovers sharper details on the head region, where specular highlights are preserved more faithfully compared to DFASR. In the second scene (Downtown West), our method produces clearer structures on the pillars and achieves more accurate color reconstruction, while also maintaining sharper edges in fine elements such as tables. The third scene (Factory) highlights the advantage of SWT in handling challenging color regions, where our model more reliably reconstructs white surfaces that NoWave tends to degrade. Overall, SWT consistently preserves edges and fine details better than the baselines, producing sharper, cleaner, and more realistic outputs across both $2\times$ (first two scenes) and $3\times$ (third scene) upscaling.
@@ 5 Conclusion
Acknowledgements
The authors would like to express their sincere gratitude to Er. Saroj Shakya for his invaluable guidance and support as a co-supervisor throughout the duration of this project. The Department of Electronics and Computer Engineering, Thapathali Campus, is gratefully acknowledged for providing the opportunity and resources to carry out this research. The authors also acknowledge Ashim Bhattarai and Ashim Tiwari for their assistance with computational resources.
Data Availability Statement
The source code for both the proposed neural network and the Unreal Engine plugin used in the experiments is publicly available at https://github.com/Prateek61/WDSS. All results in the paper can be reproduced using the provided code and instructions.